# Video Is Worth a Thousand Images: Exploring the Latest Trends in Long Video Generation

### v1

@@ 4.3. Image Input With Textual Prompt And Scenes Layout
Images provide high-quality aesthetic context for video generation. image prompts provide spatial information as well as details with different granularities like the car’s color, the carpet’s texture, and the relative position and sizes of entities, which are difficult to specify in the scene description capability of LLM. We can generate or provide key entity images like actor images as a condition for the video generation pipeline. NUWA-Infinity  [54] can process both text or image as input. In VideoStudio [61], the video generation pipeline uses foreground reference images like actors or objects like the kitchen using text-to-image diffusion models. We then feed the photos and text prompts to  [61], which uses the scene reference image, the action described in the event prompt, and the camera movement in the script as inputs. Microcinema  [128] utilized a multi-stage pipeline where photorealistic images are constructed from user prompt using high-quality text to image model like SDXL [74], or DALL-E  [80]. It first generates images from text and then takes text prompts and images for video generation. VideoBooth  [41] explored video generation from reference images and user text. It projected images into text space using CLIP  [76]. VideoDrafter  [60] works on the concept of a two-stage control pipeline where the text encoder first generates an image, and then both reference image and text prompt are used to create a video. Table LABEL:input-control-papers-table captures important milestone papers with input control mechanisms.
Table 4. Input Conditioning: Milestone Papers. ’Text Prompt’ represents 4.1, ’Text, Scene Layout’ represents 4.2 and ’Image, Text, Layout’ represents 4.3
@@
Vidgen  [86] LLM generates frames, LDM with modified attention. Text Prompt 2024
@@
User textual prompts, additional metadata for scene layouts, entity descriptions, and reference images provide a rich context to the video generation model and fine alignment between user intention and generated videos. We also need long video datasets with labels or captions to train the long video generation and input control mechanism, which will be discussed in the next section.
@@ 5. Existing Datasets
The existing datasets for long video generation can be categorized as classification datasets 5.1 and captions datasets 5.2
@@ 5.1. Classification Datasets
Classification datasets contain videos labeled with different categories, such as actions, objects, scenes, or events. UCF-101  [95] is one of the earlier popular long video classification datasets with single class labels. It includes 101 action classes with 13,320 clips from 2,500 distinct videos divided into five categories: Human-Object Interaction, Body Motion Only, Human-Human Interaction, Playing Musical Instruments, and Sports. UCF101 has limitations due to the scale and number of classes. The Kinetic dataset  [9] developed by DeepMind Google improves the scale of video datasets to 306,245 videos. DeepMind released four versions of the Kinetics dataset: Kinetics-400, Kinetics-600, Kinetics-700, and Kinetics-700-2020, with the numbers representing the number of classes. Kinetics-700-2020 was the latest version, released in 2020. The Kinetics dataset contains 306,245 distinct videos, a vast improvement compared to 2500 videos from the UCF101 dataset, and is also a single class per video. Youtube-8M [1] increased the scale of videos further by curating 8 million YouTube videos with 350k+ hours of video as a multi-label dataset as compared to single label UCF101 and Kinetics. Compared to UCF and Kinetics datasets, 8M YouTube datasets contain videos on diverse topics, including gaming, food, entertainment, health, and finance. HowTo100M  [66] scaled it further by 136 million video clips with multi labels per video sourced from 1.22M narrated instructional web videos depicting humans performing and describing over 23k different visual tasks. The granularity of video labeling can be enhanced by moving from class labels to sentences that describe scenes in the video captions dataset. The video captions datasets pair clips with sentences like in MSR-VTT  [137]; researchers sampled four frames from each video and annotated them with human-labeled sentences. This evolution of video datasets represents a general theme from single-label limited video datasets to large-scale video datasets with caption-based annotation.
@@ 5.2. Captions Datasets
MSR-VTT (MSRVideo to Text) is one of the first datasets to feature natural language video descriptions. It has 41.2 hours of videos and 200K clip-sentence pairs, covering the most comprehensive categories in earlier days of video research. MSR-VTT is illustrated in Fig. 14. WebVid-2M  [4] improved the scale of the dataset by compiling two million videos and captions generated through automatic image captioning, similar to Google’s conceptual captions  [87]. InternVid  [130] took a similar approach of using automatic captioning of images and enhanced the volume of videos further by creating a dataset of 234M video clips accompanied by detailed descriptions of a total of 4.1B words. A critical aspect of video generation is using spatial and temporal clues to create more detailed descriptions of scenes and events. InternVid and WebVid-2M lack high-quality captions with spatial and temporal context due to the algorithmic nature of caption generation and, in many cases, are often short(12-30 words). The latest dataset, like VideoInstruct-100K [64], addresses the quality of caption needed for video generation. VideoInstruct-100K took a hybrid approach to improve the quality of captions and used human-assisted and semi-automatic annotation techniques. Human workers took a subset of ActivityNet  [27] and enriched it with detailed information about spatial and temporal aspects, object relationships, reasoning, scene descriptions, and the chronological sequence of events enriches the captions. Another recent dataset focused on improving the quality of video captions is MiraData  [42]. They sampled video frames and used GPT4-V to create a ”dense caption,” which covers the main subject, movements, style, backgrounds, and cameras using structured tags.
@@
With these datasets, the next challenge is measuring the quality of generated videos, which involves assessing various aspects such as the quality of individual frame images, the smoothness of frame transitions, and the alignment between the text and the generated video discussed in 6.3. Table  LABEL:dataset_long_video catalogs milestone dataset used for long video generation.
@@ 6. Performance Measures
Video generation metrics can mainly be categorized into four categories, including Image Quality Metrics 6.1, Video Quality Metrics 6.2, Semantics Quality Metrics  6.3 and Composite Metrics 6.4.
@@ 6.1. Image Quality Metrics
Image quality metrics are essential for evaluating the quality of individual image frames produced by generative models. Among the various metrics, the Inception Score (IS) [85] is one of the most widely used to assess image-generative models. Inception Score measures the quality and diversity of generated images using a pre-trained Inception model  [99]. The Inception Score combines two key factors: image quality, defined by the clarity and accuracy of object classification, and image diversity, which refers to the variety of image classes the model can generate. A high IS indicates that the model produces diverse images that are identifiable across various categories. However, there are significant limitations to IS. While it measures the classification accuracy of generated images, it does not account for their perceptual quality—the sharpness, realism, or naturalness that a human observer might judge. Additionally, because the Inception model was trained on ImageNet, it struggles to generalize to images from domains outside of ImageNet, such as medical images, artwork, or other specialized datasets. Fréchet Inception Distance (FID) [29] was proposed to overcome the limitations of IS. FID directly measures the quality of generated images by comparing their feature distributions to those of natural images. The FID metric has been widely adopted because it provides a more comprehensive evaluation of generated image quality, as it is less reliant on specific class labels and better captures the perceptual and statistical properties of the generated images. Although image metrics can measure frame quality, they are limited in measuring flow and motion transition between frames. Video metrics attempt to solve these issues.
@@ 6.2. Video Quality Metrics
To qualitatively assess the generated videos, we need metrics that measure the visual quality and temporal coherence of the generated samples. This task is more challenging than evaluating the quality of images, as in videos, both spatial and temporal dimensions must be analyzed together. More than traditional image quality metrics like FID are required to capture the dynamics and temporal relationships unique to video generation. One such metric adapted for this purpose is the Fréchet Video Distance (FVD)  [116], which extends the FID distance to handle video data. FVD utilizes a 3D convolutional network to extract features from video clips, effectively capturing the temporal dependencies essential to video content. It allows FVD to evaluate the quality of individual frames and the consistency of the generated video’s temporal relationships (e.g., motion, transitions, and scene continuity).
However, like FID, FVD remains computationally expensive, as it involves processing large volumes of video data and requires significant computational resources. Despite its utility, FVD and FID primarily focus on measuring the distributional similarity between accurate and generated data. They do not directly assess aesthetic quality or technical flaws such as blur, noise, or flicker. It brings us to another important category of metrics designed to address these aspects. One such metric is Dover  [133], which stands for Disentangled Objective Video Quality Evaluator. Dover is intended to evaluate the aesthetic and technical quality of generated videos, including issues like motion blur, image noise, camera shaking, and flicker. Dover provides a more nuanced video quality evaluation by disentangling these technical and aesthetic components.
Dover created DIVIDE-3k dataset  [133], the first User-Generated Content Video Quality Assessment (UGC-VQA) database. DIVIDE-3k includes over 450,000 subjective quality opinions on videos from aesthetic and technical perspectives, helping researchers understand how these factors collectively influence the overall quality score. Some samples from the dataset are illustrated in Fig. 15. Motion quality is another dimension of generated videos, in addition to aesthetics. RAFT (Recurrent All-Pairs Field Transforms)  [112] captures motion patterns and temporal coherence between frames. RAFT is a method for estimating optical flow, which measures the apparent motion of objects across successive video frames. By computing the optical flow, RAFT can describe the movement and transformation of objects, which is crucial for evaluating the temporal alignment of generated videos.
@@
While video quality metrics like FVD and Dover work well for assessing the technical quality of generated videos, they have limitations in measuring how well a generated video aligns with the user’s intentions or the semantic content outlined in the prompts. To address this gap, we must explore semantic alignment metrics and composite metrics, which combine technical quality and alignment with user-defined content. These metrics are designed to evaluate how well the generated video reflects the intentions expressed in the input prompts.
@@ 6.3. Semantics Alignment Metrics
Semantic quality metrics are designed to evaluate the alignment between generated videos and the user’s intentions, specifically assessing how well the content of a video matches the descriptions or prompts provided as input. One of the most widely used tools in this domain for image-to-text alignment is CLIP (Contrastive Language-Image Pre-training)  [76]. CLIP uses contrastive learning to jointly embed text and images into a shared space, allowing it to measure the semantic similarity between text descriptions and their corresponding visual representations. By calculating the distance between the embeddings of the caption and the image, CLIP measures how well the image corresponds to the description. An extension of CLIP for assessing image-text alignment is CLIPScore  [28]. CLIPScore utilizes the embeddings generated by the pre-trained CLIP model to calculate the similarity between an image and a given textual prompt. This metric has become popular because it is reference-free and does not require ground truth or human-labeled data to compute its score. It offers an efficient way to assess the quality of image generation models based on how accurately the generated images align with the provided textual descriptions. While CLIP and CLIPScore are potent tools for measuring the semantic alignment of images and text, they focus primarily on the visual and textual features that can be easily represented as embeddings. These metrics are often limited by their inability to capture complex interactions between objects, actions, or attributes within an image. To address this, newer approaches like GRiT (Generative Region-to-Text Transformer)  [134] have been introduced to enhance the semantic understanding in generative models by going beyond simple image-text matching. GRiT is particularly notable for its ability to focus on high-level object understanding within images, allowing it to interpret not just simple nouns (e.g., ”dog,” ”car”) but also more complex and descriptive sentences that include object attributes or actions (e.g., ”a brown dog running across a field” or ”a car speeding through the city”). By incorporating such fine-grained descriptions, GRiT can provide a much richer understanding of the objects and actions in a generated image, making it more capable of aligning the generated content with the semantic intent of the user.
@@
Table 6. Metrics for video quality evaluation. In the ”Direction” column, $\uparrow$ represents higher score is better, while $\downarrow$ represents that a lower score is better.
@@
As illustrated in Fig. 17, GRiT uses a transformer-based architecture to learn the relationships between different regions of an image and their corresponding textual descriptions. It allows the model to break down the image into distinct regions and understand how each part corresponds to specific prompt components. In conclusion, while CLIP and CLIPScore provide practical ways to measure the similarity between images and text, GRiT offers a significant advancement by enabling a deeper semantic understanding of the content within images. By considering not only individual objects but also their relationships, attributes, and actions, GRiT enhances the ability to evaluate generated content on a much more complex and nuanced level. These advances in semantic quality metrics are crucial for improving the alignment of generated videos with user intentions and ensuring that the videos are visually accurate and semantically meaningful.
@@ 6.4. Composite Metrics
Composite metrics are aggregates of different individual metrics and use a suite of text prompts in various categories like animals, plants, objects, people, etc, as well as other aspects like spatial, temporal, motion, and alignment. They use algorithms and manual labeling efforts to score text-to-video models on these metrics. FETV Bench [59] is one of the first metrics-based compositions of prompts. FETV is a multi-aspect prompt creation system that categorizes prompts based on three orthogonal aspects: content, attributes, and prompt complexity. The context and control are further divided into spatial and temporal as illustrated in Fig.  18. They have done a manual evaluation on four models and also used automatic metrics like CLIPScore [28], BLIPScore [63], and OtterVQA. UGVQ  [153] (Unified Generated Video Quality) assesses generated videos’ spatial, motion, and semantic fidelity through a comprehensive, unified framework. It constructed a Large-scale Generated Video Quality assessment LGVQ  [153] dataset to evaluate the three most critical quality dimensions: spatial quality, temporal quality, and text-to-video alignment. They partition the text prompts into foreground, background, and motion. These categories have further divided these aspects into cross-cutting themes like animals, peoples, objects, and plants. They selected six text-to-video models as illustrated in Fig. 16. They used manual labeling to score these models on these aspects.
T2V-CompBench  [96] takes a similar approach as FETV Bench, and evaluates videos’ quality and semantic fidelity based on various aspects of prompt composition, including attributes, objects, spatial relationships, actions, motion binding, and numeric accuracy. The benchmark includes a dataset of 700 prompts across seven categories. It uses manual human feedback and automated metrics such as CLIPScore  [28] and BLIP-VQA [51] for evaluation but has concluded that these automatic metrics do not correlate poorly with human metrics. VBench [39] is also a suite of prompts to evaluate video quality. It used human feedback and automatic models like Dino  [7], CLIP  [76], RAFT [112] for video quality. Text alignment uses MUSIQ [45] and GRiT  [134]. MiraBench [42] also uses a similar approach. It has a prompt dataset of 150 prompts examining six perspectives, including temporal consistency, temporal motion strength, 3D consistency, visual quality, text-video alignment, and distribution consistency. These composite benchmarks have the limitation that they need manual labor to evaluate. One challenge for research is to create models that can algorithmically evaluate these metrics, i.e., developing reliable, automated metrics for long video generation remains crucial to accurately evaluate temporal consistency, semantic fidelity, and visual quality and is one of the active research areas. Table  LABEL:metrics_for_video_generation summarizes the important metrics for video evaluation.

### v2

@@ 5.3. Image Input With Textual Prompt And Scenes Layout
Images enhance video generation by providing high-quality spatial and aesthetic details (e.g., object textures, entity positions) beyond text descriptions. NUWA-Infinity  [79] supports both text and image inputs, while VideoStudio  [125] leverages reference images (e.g., actors, objects) alongside text prompts to guide generation. Microcinema  [118] employs a multi-stage pipeline, first generating images via SDXL  [126] or DALL-E  [49], then using them for video synthesis. Similarly, VideoBooth  [127] projects reference images into CLIP  [26] text space, and VideoDrafter  [119] uses a two-stage control pipeline (text-to-image, then image+text-to-video). Key works are compared in Table LABEL:input-control-papers-table.
Table 4. Input Conditioning: Milestone Papers. ’Text Prompt’ represents 5.1, ’Text, Scene Layout’ represents 5.2 and ’Image, Text, Layout’ represents 5.3
@@
User textual prompts, additional metadata for scene layouts, entity descriptions, and reference images provide a rich context for the video generation model, facilitating fine alignment between user intention and generated videos. We also require extensive video datasets with labels or captions to train the long video generation and input control mechanisms, which will be discussed in the next section.
@@ 6. Existing Datasets
The existing datasets for long video generation can be categorized as classification datasets 6.1 and captions datasets 6.2
@@ 6.1. Classification Datasets
Video classification datasets have evolved significantly in scale and annotation granularity. Early datasets like UCF-101  [129] (13,320 clips, 101 action classes) were limited to single-label categorization. The Kinetics series  [130] expanded this to 306,245 videos across 700 classes while maintaining single-label classification. YouTube-8M  [131] introduced multi-label annotation at scale (8M videos, 350k+ hours). HowTo100M  [132] further advanced this with 136M clips featuring 23k tasks and narrative descriptions. The shift toward richer annotations is exemplified by captioning datasets like MSR-VTT  [133], which pairs video frames with descriptive sentences. This progression reflects a broader trend from constrained single-label datasets to large-scale, multi-modal video-text collections.
@@ 6.2. Captions Datasets
MSR-VTT pioneered natural language video descriptions with 41.2 hours of videos and 200K clip-sentence pairs (Fig. 17). Later datasets dramatically scaled up volume through automated methods: WebVid-2M  [134] compiled 2M videos with algorithmic captions (similar to Conceptual Captions  [135]). In contrast, InternVid  [136] expanded to 234M clips with 4.1B words. However, these lack detailed spatiotemporal context due to limitations in automated captioning. Recent datasets address this quality gap: VideoInstruct-100K  [137] enriched ActivityNet  [138] subsets with human-annotated spatial/temporal details, and MiraData  [139] employed GPT4-V to generate structured ”dense captions” covering subjects, motion, and scene attributes.
@@
Here are examples from the MSR-VTT dataset showcasing video clips paired with labeled sentences. Each example includes four frames representing the video clip and five human-generated sentences that describe the content  [133]. With these datasets, the next challenge is measuring the quality of generated videos, which involves assessing various aspects, such as the quality of individual frame images, the smoothness of frame transitions, and the alignment between the text and the generated video, as discussed in 7.3. Table  LABEL:dataset_long_video catalogs milestone dataset used for long video generation.
@@ 7. Performance Measures
Video generation metrics can be mainly categorized into four categories, including Image Quality Metrics 7.1,  textit Video Quality Metrics 7.2,  textit Semantics Quality Metrics 7.3, and  textit Composite Metrics  ref sec:compositequality.
@@ 7.1. Image Quality Metrics
Image quality metrics are critical for evaluating generative models, with the Inception Score (IS)  [147] being a widely adopted measure. IS uses a pre-trained Inception model  [148] to assess quality (classification accuracy) and diversity (variety of classes) but fails to capture perceptual quality or generalize beyond ImageNet domains. To address these limitations, Fréchet Inception Distance (FID)  [149] was introduced, comparing feature distributions between generated and authentic images for a more robust evaluation of perceptual and statistical fidelity. While these metrics excel for individual frames, they cannot assess temporal dynamics, such as motion transitions, a gap addressed by video-specific metrics.
@@ 7.2. Video Quality Metrics
Video quality assessment requires metrics that evaluate both spatial and temporal coherence, extending beyond traditional image metrics such as FID. The Fréchet Video Distance (FVD)  [150] extends FID to videos by using 3D convolutions to capture temporal dynamics (motion, transitions), though it remains computationally intensive and doesn’t assess aesthetic/technical flaws. To address this, Dover  [151] evaluates technical quality (blur, noise, flicker) and aesthetics, leveraging its DIVIDE-3k dataset with 450K+ subjective annotations (Fig. 18). For motion analysis, RAFT  [152] measures optical flow to quantify object movement and temporal alignment between frames. Together, these metrics provide complementary insights into visual fidelity, temporal coherence, and motion quality.
@@
While video quality metrics like FVD and Dover work well for assessing the technical quality of generated videos, they have limitations in measuring how well a generated video aligns with the user’s intentions or the semantic content outlined in the prompts. To address this gap, we must explore semantic alignment metrics and composite metrics, which combine technical quality and alignment with user-defined content. The metrics presented in Table LABEL:video_metrics assess how accurately the generated videos align with the intentions described in their input prompts. Additionally, Table LABEL:metrics_for_video_generation highlights the evolution of video quality and semantic alignment metrics over time, showcasing advancements in generative models.
@@ 7.3. Semantics Alignment Metrics
Semantic quality metrics assess how well-generated videos align with user intent, particularly in response to text prompts. CLIP  [26] is a foundational tool for image-text alignment, using contrastive learning to embed both modalities into a shared space and measure their semantic similarity. Its extension, CLIPScore  [153], leverages CLIP’s embeddings to provide a reference-free metric for assessing image-text correspondence without ground-truth labels, making it efficient for evaluating generation models.
CLIPSIM [154] extends this paradigm for videos. CLIPSIM computes the similarity between the text and each video frame and then averages these scores to measure semantic matching. Table LABEL:tab:complete_video_metrics_v4 shows how models have been improved on these frame-level metrics and their derivatives over time.
While CLIP and CLIPScore effectively measure basic image-text alignment through embeddings, they struggle with complex object interactions or nuanced descriptions (e.g., attributes and actions). GRiT  [155] addresses this by leveraging region-to-text understanding, enabling finer-grained interpretation of scenes (e.g., “a brown dog running”) and better alignment with user intent.
@@
Table 6. Metrics for video quality evaluation. In the “Direction” column, $\uparrow$ represents higher score is better, while $\downarrow$ represents that a lower score is better.
@@
As illustrated in Fig. 20, GRiT employs a transformer-based architecture to learn the relationships between different image regions and their corresponding textual descriptions. It enables the model to break down the image into distinct regions and understand how each part corresponds to specific components of the prompt. In conclusion, while CLIP and CLIPScore provide practical methods for measuring the similarity between images and text, GRiT offers a significant advancement by enabling a deeper semantic understanding of the content within images. By considering not only individual objects but also their relationships, attributes, and actions, GRiT enhances the ability to evaluate generated content on a much more complex and nuanced level. These advances in semantic quality metrics are crucial for enhancing the alignment of videos generated with user intentions, ensuring that the videos are both visually accurate and semantically meaningful.
Table 7. Comprehensive benchmark comparison of video generation models ordered by publication year (2017-2024)  [161],  [11],  [162], [21],[101], [22], [100], [163], [164], [165], [166], [107], [128], [167], [127], [106], [108], [109], [96], [95]
Model (Citation) Year Zero-Shot Samples FVD FID CLIPSIM IS
2021 Models
GODIVA [154] 2021 No 30 0.2402
NUWA [168] 2021 X 0.97M 28.46
VideoGPT [47] 2021 103.3 24.69 24.69
Video Transformer [62] 2021 94 ± 2
TGAN-F [169] 2021 7817 22.91
LVT [170] 2021 125.8
VGAN [171] 2021 8.31
2022 Models
Make-A-Video [162] 2022 Yes 1 367.23 13.17 0.3049 33.00
CogVideo (Chinese) [14] 2022 Yes 1 701.59 23.59 0.2614 23.55
CogVideo (English) [14] 2022 Yes 1 701.59 0.2631 25.27
VideoFusion [172] 2022 639.90
LVDM [163] 2022 372.00
Video Diffusion [173] 2022 295 57
Phenaki [63] 2022 Yes 15M 37.74
MagicVideo [174] 2022 655.00
2023 Models
PYoCo [175] 2023 355.19 0.3204 47.76
VideoPoet (Pretrain) [161] 2023 355 0.3049 38.44
VideoFactory [176] 2023 410.00 0.3005
ModelScope [177] 2023 410.00 12.32 0.2930
LaViE [178] 2023 526.30 0.2949
Video LDM [179] 2023 550.61 0.2929 33.45
Vlogger [164] 2023 Yes 10M 292.43 37.23
ModelScopeT2V [177] 2023 12.32 0.2909
VideoDirectorGPT [166] 2023 12.22 0.2860
InternVid [180] 2023 617 0.2951 21.04
MicroCinema [165] 2023 342.86
PixelDance [181] 2023 242.82 42.10
Emu-Video [182] 2023 317.10 42.70
2024 Models
Lumiere [183] 2024 332.49
Reducio-DiT [109] 2024 318.50
Goku-2B (256×256)[108] 2024 Yes 246.17 45.77 ± 1.10
FVD: UCF-101 metrics (from Video LDM comparison table) FID/CLIPSIM: MSR-VTT metrics
BAIR FVD scores shown in FVD column X: Not applicable, Yes: Zero-shot capable
ModelScope replication results Bold: Best results in each category
@@ 7.4. Composite Metrics
Composite metrics combine multiple evaluation approaches to assess text-to-video generation across diverse categories (animals, objects, people) and dimensions (spatial, temporal, motion alignment). FETV Bench  [156] pioneered this approach with its three-aspect prompt system (content, attributes, complexity Fig 21) and hybrid evaluation using both manual labeling and automated metrics like CLIPScore  [153] and BLIPScore  [184]. UGVQ  [146] expanded this framework through its LGVQ dataset, evaluating spatial quality, motion coherence, and text-video alignment with partitioned prompts (foreground/background/motion) and manual scoring of six models (Fig. 19). Subsequent benchmarks, such as T2V-CompBench, VBench/VBench++, and MiraBench, have further developed comprehensive evaluation protocols, although they reveal persistent gaps in the reliability of automated metrics - particularly for assessing long video temporal consistency and semantic fidelity (Table LABEL:FETVMetrics).
Table 8. Comparative analysis of video generation models across 12 VBench dimensions [156, 157]
Models Subj. Cons. Bkg. Cons. Temp. Flicker Motion Aesthetic Obj. Class
LaVie [185] 91.41 97.47 98.30 96.38 54.94 91.82
ModelScope [186] 89.87 95.29 98.28 95.79 52.06 82.25
CogVideo [14] 92.19 96.20 97.64 96.47 38.18 73.40
VideoCrafter [187] 96.85 98.22 98.41 97.73 63.13 92.55
Gen-2 [188] 97.61 97.61 99.56 99.58 66.96 90.92
AnimateDiff [189] 95.30 97.68 98.75 97.76 67.16 90.90
Latte-1 [190] 88.88 95.40 98.89 94.63 61.59 86.53
Pika-1.0 [191] 96.94 97.36 99.74 99.50 62.04 88.72
Kling [192] 98.33 97.60 99.30 99.40 61.21 87.24
Gen-3 [188] 97.10 96.62 98.61 99.23 63.34 87.81
CogVideoX [193] 96.23 96.52 98.66 96.92 61.98 85.23
Models Mult. Obj. Human Act. Color Spatial Temp. Style Overall
LaVie [185] 33.32 96.80 86.39 34.09 25.93 26.41
ModelScope [186] 38.98 92.40 81.72 33.68 25.37 25.67
CogVideo [14] 18.11 78.20 79.57 18.24 7.80 7.70
VideoCrafter [187] 40.66 95.00 92.92 35.86 25.84 28.23
Gen-2 [185] 55.47 89.20 89.49 66.91 24.12 26.17
AnimateDiff [189] 36.88 92.60 87.47 34.60 26.03 27.04
Latte-1 [190] 34.53 90.00 85.31 41.53 24.76 27.33
Pika-1.0 [191] 43.08 86.20 90.57 61.03 24.22 25.94
Kling [192] 68.05 93.40 89.90 73.03 24.17 26.42
Gen-3 [188] 53.64 96.40 80.90 65.09 24.71 26.69
CogVideoX [193] 62.11 99.40 82.81 66.35 25.38 27.59
Table 8. Video generation model comparison (continued)